# Newton’s method for nonlinear mappings into vector bundles

### v1

@@ 7. Applications
In this section we discuss the two simplest applications of Newton’s method on vector bundles. First, the tangent bundle of a Banach manifold $\mathcal{X}$ is considered. Like in [6, 2], Newton’s method is then used to find fixed points of vector fields $\nu\in\Gamma(T\mathcal{X})$. Second, considering the cotangent bundle $T^{*}\mathcal{X}$, Newton’s method can also be applied in optimization on manifolds and to find stationary points of covector fields.
If $\mathcal{X}$ is a Riemannian manifold, in both applications the Levi-Civita connection (see e.g. [13]) is often used to compute the Newton direction. This connection induces the parallel transport along geodesics and the exponential map. Computing these classical parallel transports can be computationally very expensive. Thus, instead of starting with a connection and deriving a vector transport from it, we take the opposite route. We start with a computationally tractable vector back-transport $V^{-1}_{y}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$, and derive a consistent connection $Q$ by differentiation. Consistency is needed to apply our globalization strategy. Since retractions are anyway used in algorithms, we will discuss how vector transports on $T\mathcal{X}$ and $T^{*}\mathcal{X}$ can be derived from retractions.
@@ 7.1. Fixed points of vector fields
Consider the tangent bundle $\pi:T\mathcal{X}\to\mathcal{X}$ of a Banach manifold $\mathcal{X}$ and a vector field $\nu\in\Gamma(T\mathcal{X})$, i.e. $\nu:\mathcal{X}\to T\mathcal{X}$ with $\pi(\nu(x))=x\;\forall x\in\mathcal{X}$. Our aim is to apply Newton’s method on $\nu:\mathcal{X}\to T\mathcal{X}$ to compute a zero $x\in\mathcal{X}$ of the vector field, i.e. a point $x\in\mathcal{X}$ such that
@@
The computation of the Newton direction requires a connection map $Q:T(T\mathcal{X})\to T\mathcal{X}$, which we define with the help of a vector back-transport $V^{-1}_{x}\in\Gamma(\mathcal{L}(T\mathcal{X},T_{x}\mathcal{X}))$. As seen before in Lemma 2.3, differentiating ${\langle{V_{x}^{-1}}\rangle}\in C^{1}(T\mathcal{X},T_{x}\mathcal{X})$, leads to the consistent connection map at $\nu(x)\in T\mathcal{X}$ with $\pi(\nu(x))=x$:
@@
Vector back-transports derived from retractions.
@@
$V_{x}^{-1}(\xi):=R_{x}^{\prime}(R_{x}^{-1}(\xi))^{-1}:T_{\xi}\mathcal{X}\to T_{x}\mathcal{X},$
@@
In the Newton iteration we use $\xi=x_{+}=R_{x}(\delta x)$, and thus $R_{x}^{-1}(\xi)=\delta x$. Hence, $V_{x}^{-1}(\xi)=R_{x}^{\prime}(\delta x)$, and the inverse retraction is not needed for the evaluation of the vector back-transport.
@@
Embedded submanifolds.
If $\mathcal{X}$ is an embedded submanifold of a Hilbert space $H$ we can use the embedding $E_{H}:T\mathcal{X}\to H$ with $E_{H}(\xi):T_{\xi}\mathcal{X}\to H$ and the orthogonal projection $P(x):H\to T_{x}\mathcal{X}$ on the tangent space to define a vector back-transport locally by
$V_{x}^{-1}(\xi):=P(x)E_{H}(\xi)\in L(T_{\xi}\mathcal{X},T_{x}\mathcal{X})$
Consider a vector field $\nu:\mathcal{X}\to T\mathcal{X}$. By our natural inclusion $\varphi$, we obtain the $C^{1}$-mapping ${\langle{V_{x}^{-1}}\rangle}:T\mathcal{X}\to T_{x}\mathcal{X}$ with ${\langle{V_{x}^{-1}}\rangle}(\nu)(\xi)=P(x)(E_{H}\nu)(\xi)$. Differentiating this with respect to $\xi$ at $\xi=x$, we obtain
$Q_{\nu(x)}\nu^{\prime}(x)\delta x={\langle{V_{x}^{-1}}\rangle}^{\prime}(\nu(x))\nu^{\prime}(x)\delta x=P(x)(E_{H}\nu)^{\prime}(x)\delta x\in T_{x}\mathcal{X},$
which defines the consistent connection map $Q_{\nu(x)}={\langle{V_{x}^{-1}}\rangle}^{\prime}(\nu(x))$. The covariant derivative $Q\circ\nu^{\prime}$ of a vector field $\nu$ w.r.t. the connection map $Q$ can therefore be determined by projecting the euclidean directional derivatives onto $T\mathcal{X}$ by the orthogonal projection. The connection which belongs to this covariant derivative is sometimes called tangential connection. It actually defines the Levi-Civita connection on $T\mathcal{X}$, see e.g. [14, Chap. 5].
@@ 7.2. Variational problems on manifolds
Consider a twice differentiable function $f:\mathcal{X}\rightarrow\mathbb{R}$, defined on a Banach manifold $\mathcal{X}$. In order to find a critical point, we want to apply Newton’s method on the cotangent bundle $T^{*}\mathcal{X}$ to compute a zero of the covector field $f^{\prime}:\mathcal{X}\rightarrow T^{*}\mathcal{X}$, i.e. a point $x\in\mathcal{X}$ such that
$f^{\prime}(x)=0_{x}^{*}\in T_{x}\mathcal{X}^{*}.$
Thus, we need a dual connection map $Q_{f^{\prime}(x)}^{*}:T_{f^{\prime}(x)}(T^{*}\mathcal{X})\rightarrow T_{x}^{*}\mathcal{X}$ on the dual bundle $T^{*}\mathcal{X}$. Again, such a connection map can be derived from a vector transport. We will first discuss dual connection maps in general and come back to the cotangent bundle afterwards.
@@
Let $p:\mathcal{E}\to\mathcal{M}$ be a vector bundle. The dual bundle $p^{*}:\mathcal{E}^{*}\to\mathcal{M}$ is defined as a vector bundle with fibres via $(p^{*})^{-1}(y)=E_{y}^{*}$, i.e. the dual space of $E_{y}$. We can view this as a special case $\mathcal{E}^{*}=\mathcal{L}(\mathcal{E},\mathcal{M}\times\mathbb{R})$. For example the dual bundle of the tangent bundle $T\mathcal{X}$ is the cotangent bundle $T^{*}\mathcal{X}$.
We denote the dual pairing for $\ell\in\Gamma(\mathcal{E}^{*})$ and $e\in\Gamma(\mathcal{E})$ by $\ell(e):\mathcal{M}\rightarrow\mathbb{R}$. Its derivative at a point $y\in\mathcal{M}$ is given by $\ell(e)^{\prime}(y):T_{y}\mathcal{M}\to\mathbb{R}$. If $\mathcal{E}$ admits a connection map it is possible to define a dual connection map on the dual bundle $\mathcal{E}^{*}$ as follows [14, Chap. 4].
@@
Let $Q$ be a connection map on the vector bundle $p:\mathcal{E}\rightarrow\mathcal{M}$. The dual connection map $Q^{*}:T\mathcal{E}^{*}\to\mathcal{E}^{*}$ on the dual vector bundle $p^{*}:\mathcal{E}^{*}\rightarrow\mathcal{M}$ is defined by
$((Q^{*}\circ\ell^{\prime})(y)\delta y)(e):=\ell(e)^{\prime}(y)\delta y-\ell((Q\circ e^{\prime})(y)\delta y)\;\forall e\in\Gamma(\mathcal{E}),\;\forall\delta y\in T_{y}\mathcal{M}$
@@
Thus, $(Q^{*}\circ\ell^{\prime})(y)\delta y)(e)$ does not depend on $e_{\tau}^{\prime}(y)$, but on $e_{\tau}(y)$ only, so indeed $(Q^{*}\circ\ell^{\prime})(y)\delta y\in E_{y}^{*}$ with the representation
$Q^{*}_{\ell}\delta\ell\sim Q^{*}_{\ell,\tau}(\delta y,\delta\ell_{\tau})=\delta\ell_{\tau}(\cdot)+\ell\circ B_{y,\tau}(\cdot)\delta y\in\mathbb{E}^{*}.$ (28)
Comparison of (27) and (28) shows that if the dual connection map $Q^{*}$ corresponds to a connection map $Q$ on $\mathcal{E}$ the bilinear form $B_{y,\tau}$ is used in both representations.
Now consider a vector transport $V_{y}\in\Gamma(\mathcal{L}(\mathcal{M}\times\mathcal{E},E_{y})$. We can define a covector back-transport $V_{y}^{*}\in\Gamma(\mathcal{L}(\mathcal{E}^{*},E_{y}^{*}))$, which acts as a vector back-transport on the dual bundle by using pointwise adjoints $V_{y}^{*}(\hat{y}):=(V_{y}(\hat{y}))^{*}\in L(E_{\hat{y}}^{*},E_{y}^{*})$, i.e.
@@
The choice (29) defines the dual connection map on $\mathcal{E}^{*}$ corresponding to the connection map $Q_{e}={\langle{V_{y}^{-1}}\rangle}^{\prime}(e)$ at $e$ with $y=p(e)=p^{*}(\ell)$ on $\mathcal{E}$ derived from the vector transport $V_{y}$.
@@
Let $\ell\in\Gamma(\mathcal{E}^{*})$, $e\in\Gamma(\mathcal{E})$, and $\delta y\in T_{y}\mathcal{M}$. In trivializations, if $\ell(y)$ is represented by $(y,\ell_{\tau}(y))$ and $\ell^{\prime}(y)\delta y$ by $(\delta y,\delta\ell_{\tau})$, we obtain by using ${\langle{V_{y}^{*}}\rangle}(\ell)=\ell\circ V_{y}(p^{*}(\ell))$ and (2):
@@
Since $B_{y,\tau}(e_{\tau}(y))\delta y=-((V_{y,\tau}^{-1})^{\prime}(y)\delta y)e_{\tau}(y)=(V_{y,\tau}^{\prime}(y)\delta y)e_{\tau}(y)$ by (4) and (5), we obtain the desired result. ∎
Remark 7.4.
$\nabla_{\delta y}e(y):=Q\circ e^{\prime}(y)\delta y\mbox{ for }y\in\mathcal{M},e\in\Gamma(\mathcal{E}),\delta y\in\Gamma(T\mathcal{M}),$
in which case Definition 7.2 yields a product rule:
On Riemannian manifolds the Levi-Civita connection is widely used on $T\mathcal{M}$. In this case, dual connections can also be defined with the help of a Riesz isomorphism $\mathcal{R}:T\mathcal{M}\to T^{*}\mathcal{M}$:
$\nabla_{\delta y}^{*}\ell:=\mathcal{R}(\nabla_{\delta y}\mathcal{R}^{-1}(\ell)).$
However, this expression is often expensive to evaluate numerically.
Covector back-transports via retractions.
Consider the cotangent bundle $\pi^{*}:T^{*}\mathcal{X}\to\mathcal{X}$. Since the cotangent bundle is the dual vector bundle to the tangent bundle it admits a corresponding dual connection map. In order to define the dual connection map which fits to the connection map on $T\mathcal{X}$ induced by the retraction, we first define the covector back-transport corresponding to $V_{x}\in\Gamma(\mathcal{X}\times T_{x}\mathcal{X},T\mathcal{X})$ given by (25). For $v^{*}\in T_{\xi}^{*}\mathcal{X}$ we set
$V^{*}_{x}(\xi)(v^{*}):=v^{*}\circ R_{x}^{\prime}(R_{x}^{-1}(\xi))=R_{x}^{\prime}(R_{x}^{-1}(\xi))^{*}(v^{*})\in T_{x}^{*}\mathcal{X}.$

### v2

@@ 7. Applications
In this section we discuss some applications of Newton’s method on vector bundles, and we provide ways to construct the required connections, with the help of available algorithmic quantities. As a first example, the tangent bundle of a Banach manifold $\mathcal{X}$ is considered, where, like in [6, 2], Newton’s method is used to find fixed points of vector fields. Second, considering the dual bundle $\mathcal{E}^{*}$ of a vector bundle $\mathcal{E}$, we study, how Newton’s method can be applied to variational problems on manifolds. In particular, we will discuss the application of Newton’s method on covector fields to find stationary points of functionals and relate our method to Newton-SQP-methods on manifolds. In the last section of this chapter we will consider the important special case of embedded vector bundles.
@@ 7.1. Fixed points of vector fields
Consider the tangent bundle $\pi:T\mathcal{X}\to\mathcal{X}$ of a $C^{2}$-Banach manifold $\mathcal{X}$ and a vector field $\nu\in\Gamma(T\mathcal{X})$, i.e. $\nu:\mathcal{X}\to T\mathcal{X}$ with $\pi(\nu(x))=x$ for all $x\in\mathcal{X}$. Our aim is to apply Newton’s method on $\nu:\mathcal{X}\to T\mathcal{X}$ to compute a zero $x\in\mathcal{X}$ of the vector field, i.e. a point $x\in\mathcal{X}$ such that
@@
The computation of the Newton direction requires a connection map $Q\in\Gamma(\mathcal{L}(T(T\mathcal{X}),T\mathcal{X}))$, which we define with the help of a vector back-transport $V^{-1}_{x}\in\Gamma(\mathcal{L}(T\mathcal{X},T_{x}\mathcal{X}))$. As seen before in Lemma 2.3, differentiating ${\langle{V_{x}^{-1}}\rangle}\in C^{1}(T\mathcal{X},T_{x}\mathcal{X})$, leads to the consistent connection map at $\nu(x)\in T\mathcal{X}$ with $\pi(\nu(x))=x$:
@@
Vector back-transports via retractions.
@@
$V_{x}^{-1}(\xi):=R_{x}^{\prime}(R_{x}^{-1}(\xi))^{-1}:T_{\xi}\mathcal{X}\to T_{x}\mathcal{X}.$ (35)
@@
In the Newton iteration we use $\xi=x_{+}=R_{x}(\delta x)$, and thus $R_{x}^{-1}(\xi)=\delta x$. Hence, $V_{x}^{-1}(\xi)=R_{x}^{\prime}(\delta x)^{-1}$, and the inverse retraction is not needed for the evaluation of the vector back-transport.
@@ 7.2. Variational problems on manifolds
In mathematical physics we frequently encounter variational problems, which often yield differential equations in weak form. If these problems are posed on manifolds (as, for example, in variational mechanics [17], harmonic maps [10] or Cosserat materials [22]), dual bundles can be used for their treatment. We thus will consider Newton’s method for a mapping $F\in C^{1}(\mathcal{X},\mathcal{E}^{*})$ from a manifold $\mathcal{X}$ into a dual bundle $\mathcal{E}^{*}$.
Let $\mathcal{X}$ be a Banach manifold, $p:\mathcal{E}\to\mathcal{Y}$ a vector bundle and $p^{*}:\mathcal{E}^{*}\to\mathcal{Y}$ its dual bundle with fibres $E^{*}_{y}=(p^{*})^{-1}(y)=L(E_{y},\mathbb{R})$, being the dual space of $E_{y}$, so $\mathcal{E}^{*}=\mathcal{L}(\mathcal{E},\mathbb{R})$. For example the dual bundle of the tangent bundle $T\mathcal{X}$ is the cotangent bundle $T^{*}\mathcal{X}$.
If $\tau_{\mathcal{E}}:\mathcal{E}\to\mathbb{E}$ is a trivialization with $e_{\tau}=\tau_{y,\mathcal{E}}(e)$ for $e\in E_{y}$, then by (1) dispensing with a trivialization of the co-domain $\mathbb{R}$ of $\ell\in E_{y}^{*}=L(E_{y},\mathbb{R})$, we obtain $\tau_{\mathcal{E}^{*}}:\mathcal{E}^{*}\to\mathbb{E}^{*}$ by the mapping $\ell_{\tau}=\tau_{\mathcal{E}^{*}}(\ell):=\ell\circ\tau_{\mathcal{E},y}^{-1}$, so that
$\ell_{\tau}(e_{\tau})=\ell(e).$ (37)
We denote the fibrewise dual pairing for $\ell\in\Gamma(\mathcal{E}^{*})$ and $e\in\Gamma(\mathcal{E})$ by $\ell(e):\mathcal{Y}\rightarrow\mathbb{R}$. Its derivative at a point $y\in\mathcal{Y}$ is given by $\ell(e)^{\prime}(y)\in L(T_{y}\mathcal{Y},\mathbb{R})$.
For a mapping $F:\mathcal{X}\rightarrow\mathcal{E}^{*}$ with $y(x)=p^{*}(F(x))$, we would like to find the solution of the following variational problem, where $e\in E_{y(x)}$ plays the role of a test function:
$F(x)=0_{y(x)}^{*}\quad\Leftrightarrow\quad F(x)(e)=0\quad\forall e\in E_{y(x)}.$
@@
To define Newton steps on dual bundles, we need a dual connection map $Q_{\ell}^{*}:T_{\ell}(\mathcal{E}^{*})\rightarrow E^{*}_{p^{*}(\ell)}$ for $\ell\in\mathcal{E}^{*}$ and apply it to $F^{\prime}(x)\delta x\in T_{F(x)}\mathcal{E}^{*}$:
$Q_{F(x)}^{*}\circ F^{\prime}(x)\delta x+F(x)=0^{*}_{y(x)}\quad\Leftrightarrow\quad(Q_{F(x)}^{*}\circ F^{\prime}(x)\delta x+F(x))(e)=0\quad\forall e\in E_{y(x)}.$
Given a connection $Q$ on $\mathcal{E}$, $Q^{*}$ can be defined as follows [15, Chap. 4]:
@@
Let $Q$ be a connection map on $p:\mathcal{E}\rightarrow\mathcal{Y}$. The dual connection map $Q^{*}\in\Gamma(\mathcal{L}(T\mathcal{E}^{*},(p^{*})^{*}\mathcal{E}^{*}))$ on $p^{*}:\mathcal{E}^{*}\rightarrow\mathcal{Y}$ is defined by
$(Q_{\ell(y)}^{*}\circ\ell^{\prime}(y)\delta y)(e(y)):=\ell(e)^{\prime}(y)\delta y-\ell(y)(Q_{e(y)}\circ e^{\prime}(y)\delta y)\;\forall e\in\Gamma(\mathcal{E}),\;\forall\delta y\in T_{y}\mathcal{Y}.$ (38)
@@
Thus, $((Q^{*}\circ\ell^{\prime})(y)\delta y)(e)$ does not depend on $e_{\tau}^{\prime}(y)$, but on $e_{\tau}(y)$ only, so indeed $(Q^{*}\circ\ell^{\prime})(y)\delta y\in E_{y}^{*}$ with the representation for any $\delta\ell\in T_{\ell}\mathcal{E}^{*}\sim(\delta y,\delta\ell_{\tau})\in T_{y}\mathcal{Y}\times\mathbb{E}^{*}$:
$Q^{*}_{\ell}\delta\ell\in E_{y}^{*}\sim Q^{*}_{\ell,\tau}(\delta y,\delta\ell_{\tau})=\delta\ell_{\tau}(\cdot)+\ell_{\tau}(y)\circ B_{y,\tau}(\cdot)\delta y\in\mathbb{E}^{*},$ (40)
which can also be written as an equality, with $e_{\tau}=\tau_{\mathcal{E},y}(e)$:
$(Q^{*}_{\ell}\delta\ell)(e)=Q^{*}_{\ell,\tau}(\delta y,\delta\ell_{\tau})(e_{\tau})=\delta\ell_{\tau}(e_{\tau})+\ell_{\tau}(y)\circ B_{y,\tau}(e_{\tau})\delta y.$ (41)
Remark 7.3.
$\nabla_{\delta y}e(y):=Q\circ e^{\prime}(y)\delta y\mbox{ for }y\in\mathcal{Y},e\in\Gamma(\mathcal{E}),\delta y\in T_{y}\mathcal{Y},$
in which case Definition 7.2 yields a product rule:
On Riemannian manifolds the Levi-Civita connection is widely used on $T\mathcal{Y}$. In this case, dual connections can also be defined with the help of a Riesz isomorphism $\mathcal{R}:\Gamma(\mathcal{L}(T\mathcal{Y},T^{*}\mathcal{Y}))$:
$\nabla_{\delta y}^{*}\ell(y):=\mathcal{R}_{y}(\nabla_{\delta y}\mathcal{R}_{y}^{-1}(\ell(y))).$
However, this expression is often expensive to evaluate numerically, since it requires the application of $\mathcal{R}_{y}$ and $\mathcal{R}_{y}^{-1}$.
In Newton’s method we set $\ell=F(x)$ and obtain by application of (41) with representations $F^{\prime}(x)\delta x\sim(y^{\prime}(x)\delta x,(F^{\prime}(x)\delta x)_{\tau})$:
$(Q_{F(x)}^{*}\circ F^{\prime}(x)\delta x)(e)=(F^{\prime}(x)\delta x)_{\tau}(e_{\tau})+F_{\tau}(x)\circ B_{y(x),\tau}(e_{\tau})y^{\prime}(x)\delta x\quad\forall e_{\tau}=\tau_{\mathcal{E},y}(e)\in\mathbb{E}.$ (42)
Covector back-transports.
Now consider a vector transport $V_{y}\in\Gamma(\mathcal{L}(\mathcal{Y}\times\mathcal{E},E_{y}))$. We can define a covector back-transport $V_{y}^{*}\in\Gamma(\mathcal{L}(\mathcal{E}^{*},E_{y}^{*}))$, which acts as a vector back-transport on the dual bundle by using fibrewise adjoints $V_{y}^{*}(\hat{y}):=(V_{y}(\hat{y}))^{*}\in L(E_{\hat{y}}^{*},E_{y}^{*})$, i.e.
@@
In particular, the covector back-transport, needed in the simplified Newton step (33) can be performed via a forward transport of the test-vectors:
$V_{y(x)}^{*}(y(x_{+}))F(x_{+})(e)=F(x_{+})(V_{y(x)}(y(x_{+}))e)\quad\forall e\in E_{y}.$
@@
The choice (43) defines the dual connection map $Q_{\ell}^{*}$ on $\mathcal{E}^{*}$ corresponding to the connection map $Q_{e}={\langle{V_{y}^{-1}}\rangle}^{\prime}(e)$ at $e$ with $y=p(e)=p^{*}(\ell)$ on $\mathcal{E}$.
With this dual connection we have for $F\in C^{1}(\mathcal{X},\mathcal{E}^{*})$:
$(Q^{*}_{F(x)}\circ F^{\prime}(x)\delta x)(e)=\frac{d}{d\xi}\Big{(}F(\xi)(V_{y(x)}(y(\xi))e)\Big{)}\Big{|}_{\xi=x}\delta x\qquad\forall\delta x\in T_{x}\mathcal{X}.$ (44)
@@
Let $\ell\in\Gamma(\mathcal{E}^{*})$, $e\in\Gamma(\mathcal{E})$, and $\delta y\in T_{y}\mathcal{Y}$. In trivializations, if $\ell(y)$ is represented by $(y,\ell_{\tau}(y))$ and $\ell^{\prime}(y)\delta y$ by $(\delta y,\delta\ell_{\tau})$, we obtain by using ${\langle{V_{y}^{*}}\rangle}(\ell)=\ell\circ V_{y}(p^{*}(\ell))$ and (2):
@@
Since $B_{y,\tau}(e_{\tau}(y))\delta y=-((V_{y,\tau}^{-1})^{\prime}(y)\delta y)e_{\tau}(y)=(V_{y,\tau}^{\prime}(y)\delta y)e_{\tau}(y)$ by (4) and (5), we obtain the desired result.
Concerning (44) we compute, using local trivializations of $\mathcal{E}$, (37), the product rule and (42) with $B_{y(x),\tau}(e_{\tau})y^{\prime}(x)\delta x=V^{\prime}_{y(x),\tau}(y(x))y^{\prime}(x)\delta x)(e_{\tau})$:
$\frac{d}{d\xi}\Big{(}F(\xi)(V_{y(x)}(y(\xi))e)\Big{)}\Big{|}_{\xi=x}\delta x$ $=\frac{d}{d\xi}\Big{(}F_{\tau}(\xi)(V_{y(x),\tau}(y(\xi))e_{\tau})\Big{)}\Big{|}_{\xi=x}\delta x$
$=(F^{\prime}(x)\delta x)_{\tau}(e_{\tau})+(F_{\tau}(x)\circ V^{\prime}_{y(x),\tau}(y(x))y^{\prime}(x)\delta x)(e_{\tau})$
$=Q^{*}_{F(x),\tau}(y^{\prime}(x)\delta x,(F^{\prime}(x)\delta x)_{\tau})(e_{\tau})$
$=(Q^{*}_{F(x)}\circ F^{\prime}(x)\delta x)(e).$
7.3. Critical points of functionals
Consider a twice differentiable function $f:\mathcal{X}\rightarrow\mathbb{R}$, defined on a $C^{2}$-Banach manifold $\mathcal{X}$. In order to find a critical point, we want to apply Newton’s method on the cotangent bundle $\pi^{*}:T^{*}\mathcal{X}\to\mathcal{X}$ to compute a zero of the covector field $f^{\prime}:\mathcal{X}\rightarrow T^{*}\mathcal{X}$, i.e. a point $x\in\mathcal{X}$ such that
$f^{\prime}(x)=0_{x}^{*}\in T_{x}\mathcal{X}^{*}.$
Thus, we need a dual connection map $Q_{f^{\prime}(x)}^{*}:T_{f^{\prime}(x)}(T^{*}\mathcal{X})\rightarrow T_{x}^{*}\mathcal{X}$ on the cotangent bundle $T^{*}\mathcal{X}$.
Covector back-transports and dual connections via retractions.
Since the cotangent bundle is the dual vector bundle to the tangent bundle it admits a corresponding dual connection map. To define the dual connection map which fits to the connection map on $T\mathcal{X}$ induced by the retraction, we first define the covector back-transport corresponding to $V_{x}\in\Gamma(\mathcal{X}\times T_{x}\mathcal{X},T\mathcal{X})$ given by (34). For $\ell\in T_{\xi}^{*}\mathcal{X}$ we set
$V^{*}_{x}(\xi)(\ell):=\ell\circ R_{x}^{\prime}(R_{x}^{-1}(\xi))=R_{x}^{\prime}(R_{x}^{-1}(\xi))^{*}(\ell)\in T_{x}^{*}\mathcal{X}.$